OpenStudy (anonymous):

What is the volumetric flow rate (Q) in cubic feet per minute (cfm) of water traveling at 10 ft/s (v) through a 100 cm diameter pipe if Q=v*A?

OpenStudy (anonymous):

Hmm the question provided you with the formula so just use it. Q= v*A. the key here is to find A, since v is given to you already. 100 cm is equals to 3.28 ft. remember to convert if not you will not get the right answer. 3.28ft in diameter means 1.64ft in radius To find area, A= pi*r^2 A= 8.5 ft sqaure. Q= 8.5 *10 = 85 cubic feet per second. BUT the question wants cubic feet per minute. So just take 85 * 60= 510. 510 cubic feet per minute.